OpenStudy (anonymous):

A scale drawing of a game room is shown below: A rectangle is shown. The length of the rectangle is labeled 2 inches. The width of the rectangle is labeled 4.5 inches. The scale is 1 to 30. What is the area of the actual game room in square feet? Round your answer to the nearest whole number. 9 ft2 23 ft2 56 ft2 270 ft2

OpenStudy (ybarrap):

 For every 1 inch, reality is 30 inches So multiply length by 30 then multiply width by 30 then multiply both because area equals length times width 1 sq inch = 30 sq inches 2 sq inches = 2*30 sq inches ... 4 sq inches = 4*30 sq inches, etc You then convert to square inches to feet: 12 inches = 1 ft So, 12^2=144 sq inches = 1^1=1 sq ft Now divide your final answer (area in sq inches) by 144 to convert from sq inches to sq feet Area = (length*30 times width*30)/144